# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

### v4

@@ 4.2. Overall Performance (RQ1)
Higher-order CrossNetv1 (Wang et al., 2017) 0.3779 78.39 0.4456 80.59
@@
Since DCNv3 models feature interactions explicitly, we select 11 representative models for comparison, categorized into First-order, Second-order, and Higher-order classes. We bold the best performance, while underlined scores are the second best. The experimental results are shown in Table 3, and we can draw the following conclusions:
@@
  • Overall, capturing higher-order feature interactions often enhances model performance. For example, FM outperforms LR on the two large-scale datasets Avazu and Criteo, and CrossNetv2 outperforms all first-order and second-order feature interaction models except for FwFM on Avazu. This demonstrates the effectiveness of higher-order feature interactions in improving model performance.
@@
  • Overall, SDCNv3 achieves the best performance across all six datasets. SDCNv3 shows an average AUC improvement of 0.22% over the strongest baseline model and an average Logloss decrease of 0.11%, both exceeding a statistically significant performance improvement of 0.1%. This demonstrates the effectiveness of SDCNv3. SDCNv3 ranks 1st in CTR prediction on the PapersWithCode benchmarks [^9] https://paperswithcode.com/task/click-through-rate-prediction for the Criteo, KDD12, and KKBox datasets.
  • The FinalMLP model achieves good performance on the Avazu and Criteo datasets, surpassing most CTR models that combine explicit and implicit feature interactions. This demonstrates the effectiveness of implicit feature interactions. Consequently, most CTR models attempt to integrate DNN into explicit feature interaction models to enhance performance. However, SDCNv3 achieves state-of-the-art performance using only explicit feature interactions, indicating the effectiveness of modeling with explicit feature interactions alone.
@@ 4.3.1. Efficiency Comparison (RQ2)
  • The LR model has the lowest runtime at 150 seconds among all models. CrossNetv1 and CrossNetv2 follow closely, requiring only a negligible increase in time while significantly improving performance. This demonstrates the efficiency of CrossNet and its series models. As a fundamental component of deep CTR models, DNN requires only 156 seconds. Due to the parallel-friendly nature of CTR parallel structures, some carefully designed deep CTR models, such as PNN, DCNv2, FinalMLP, and FINAL, significantly enhance prediction accuracy without a substantial increase in runtime.
@@ 4.3.2. The Interpretability and noise filtering capability of SDCNv3 (RQ3)
Interpretability is a crucial aspect of CTR prediction tasks (Li et al., 2020; Wang et al., 2023a, 2021b; Song et al., 2019), as it helps researchers understand the reasons behind predictions and increases confidence in the results. In this section, we investigate the dynamic Cross & Masked Vector and the static $\mathbf{W}_{l}$ to understand the model’s prediction process. The experimental and visualization results are shown in Figure 6 (Value denotes the Frobenius norm for each feature field), we can derive the following observations:
@@
Figure 6. Visualization of field-wise value and sparsity for $\mathbf{c}_{l}$, $\texttt{Mask}(\mathbf{c}_{l})$, and $\mathbf{W}_{3}$ of SDCNv3 on a batch of ML-1M.
@@
  • $\mathbf{c}_{l}$ and $\texttt{Mask}(\mathbf{c}_{l})$ exhibit complementary properties. When a feature field in $\mathbf{c}_{l}$ becomes important, its corresponding sparsity in $\texttt{Mask}(\mathbf{c}_{l})$ decreases, and vice versa (e.g., UserID, Occupation, Age fields). This demonstrates the effectiveness of our introduced Self-Mask operation, which further filters out noise by more aggressively assigning zero values to the representation elements of certain feature fields.
@@
  • Overall, we observe that the importance of higher-order feature interactions is lower than that of lower-order feature interactions, which is similarly reported in some works (Liu et al., 2020; Wang et al., 2023a; Zhu et al., 2023a). For example, in Figure 6, (f) has fewer bright red blocks compared to (e), and the blue blocks in (a) gradually darken as the number of layers increases, the situation is similar in (b).

### v5

@@ 4.2. Overall Performance (RQ1)
High-order CrossNetv1 (Wang et al., 2017) 0.3779 78.39 0.4456 80.59
@@
Since DCNv3 models feature interactions explicitly, we select 11 representative models for comparison, categorized into First-order, Second-order, and High-order classes. We bold the best performance, while underlined scores are the second best. The experimental results are shown in Table 3, and we can draw the following conclusions:
@@
  • Overall, capturing high-order feature interactions often enhances model performance. For example, FM outperforms LR on the two large-scale datasets Avazu and Criteo, and CrossNetv2 outperforms all first-order and second-order feature interaction models except for FwFM on Avazu. This demonstrates the effectiveness of high-order feature interactions.
@@
  • Overall, SDCNv3 achieves the best performance across all six datasets, with an average AUC improvement of 0.21% over the strongest baseline model and an average Logloss decrease of 0.11%, both exceeding the statistically significant threshold of 0.1%. This demonstrates the effectiveness of SDCNv3.
  • The FinalMLP model achieves good performance on the Avazu and Criteo datasets, surpassing most CTR models that combine explicit and implicit feature interactions. This demonstrates the effectiveness of implicit feature interactions. Consequently, most CTR models attempt to integrate DNN into explicit feature interaction models to enhance performance. However, SDCNv3 achieves SOTA performance using only explicit feature interactions, indicating the effectiveness of modeling with explicit feature interactions alone.
@@ 4.3.1. Efficiency Comparison (RQ2)
  • The LR model has the lowest runtime at 150 seconds among all models. CrossNetv1 and CrossNetv2 follow closely, requiring only a negligible increase in time while improving performance. This demonstrates the efficiency of CrossNet and its series models. As a fundamental component of deep CTR models, DNN requires only 156 seconds. Due to the parallel-friendly nature of CTR parallel structures, some carefully designed deep CTR models, such as DCNv2, FinalMLP, and FINAL, significantly enhance prediction accuracy without a substantial increase in runtime.
@@ 4.3.2. The Interpretability and noise filtering capability of SDCNv3 (RQ3)
Interpretability is crucial in CTR prediction tasks (Li et al., 2020; Wang et al., 2023a, 2021b; Song et al., 2019) as it helps researchers understand predictions and increases confidence in the results. In this section, we investigate the dynamic Cross & Masked Vector and the static $\mathbf{W}_{l}$ to understand the model’s prediction process. The experimental and visualization results are shown in Figure 6 (Value denotes the Frobenius norm for each feature field), we can derive the following observations:
@@
Figure 6. Visualization of field-wise value and sparsity for $\mathbf{c}_{l}$, $\texttt{Mask}(\mathbf{c}_{l})$, and $\mathbf{W}_{3}$ of SDCNv3 on a random batch of ML-1M.
@@
  • $\mathbf{c}_{l}$ and $\texttt{Mask}(\mathbf{c}_{l})$ exhibit complementary properties. When a feature field in $\mathbf{c}_{l}$ becomes important, its corresponding sparsity in $\texttt{Mask}(\mathbf{c}_{l})$ decreases, and vice versa (e.g., UserID, Occupation, Age fields). This demonstrates the effectiveness of our introduced Self-Mask operation, which further filters out noise by more aggressively assigning zero values to certain feature fields.
@@
  • Overall, we observe that the importance of high-order feature interactions is lower than that of lower-order feature interactions, which is similarly reported in some works (Liu et al., 2020; Wang et al., 2023a; Zhu et al., 2023a). For example, in Figure 6, (f) has fewer bright red blocks compared to (e), and the blue blocks in (a) gradually darken as the number of layers increases, the situation is similar in (b).